Please help me. I an strugling with this question.
In a bumper test, three types of autos were deliberately crashed into a barrier at 5 mph, and the
resulting damage (in dollars) was estimated. Five test vehicles of each type were crashed, with
the results shown below. Research question: Are the mean crash damages the same for these threevehicles?
Crash Damage ($)
Goliath Varmint Weasel
1,600 1,290 1,090 760 1,400 2,100
880 1,390 1,830 1,950 1,850 1,250
1,220 950 1,920

Answer by checkley71(8403)

1600+760+880+1950+1220=6410/5=1282
1290+1400+1390+1850+950=880/5=1376
1090+2100+1830+1250+1920=8190/5=1638
OBVIOUSLY THE ANSWER IS NO. THESE MEAN AVERAGES ARE NOT THE SAME FOR THESE 3 VEHICLES.